OpenStudy (anonymous):

If f(x)=3+5/x, which of the following CANNOT be value of f(x)? -5/3 -3/5 0 5/3 3

OpenStudy (anonymous):

so wich one when put on y,gives x as 0

OpenStudy (anonymous):

remember hyperbola horizontal asymptote

OpenStudy (anonymous):

0, because you can't divide by 0.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that will be correct if they where looking for x-value

OpenStudy (anonymous):

y=3 is the HA-horizontal asymtote

OpenStudy (cwrw238):

if 5/x = -3 then f(x) = 0 so its not this option

OpenStudy (cwrw238):

try equating eg 3 + 5/x = -5/3 5/x = - -14/3 -14x = 15 - so it can't be first choice -5/3

OpenStudy (cwrw238):

3 is correct choice 3 + 5/x = 3 5/x = 0 x x = 5/0 which is indeterminate

OpenStudy (anonymous):

x=15/-14 is a legitimate answer, so -5/3 can be a value of x.

OpenStudy (cwrw238):

* x = 5/0

OpenStudy (cwrw238):

yes -5/3 can be a value of f(x)

OpenStudy (anonymous):

@cwrw238 u can never divide by 0 though

OpenStudy (cwrw238):

division by 0 can never be done 5 /0 is meaningless - no matter how many 0;s you add together the sum will be zero - never 5

OpenStudy (cwrw238):

correct answer is 3 f(x) = 3 + 5/x can never have a value 3